Assume that the radius of the earth is 3.00 x 103 mi. How many cubic meters of space does the earth occupy?

 



1
Expert's answer
2021-08-29T16:56:33-0400

Given:

"R=3.00*10^3{\:\rm mi}=3.00*10^3*1609{\:\rm m}\\=4.83*10^6\:\rm m"


The volume of Earth

"V=\frac{4}{3}\pi R^3"

"V=\frac{4}{3}\pi (4.83*10^6\:\rm m)^3=4.71*10^{20}\: m^3"